OpenStudy (anonymous):

y<= x - 5 y>= -x - 4

OpenStudy (allank):

hello. please clarify what the question is. what are we required to do with the inequalities?

OpenStudy (anonymous):

its graphing and theanswer choices are (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (allank):

are you using a graphing calculator?

OpenStudy (anonymous):

im not

OpenStudy (allank):

just a sec...

OpenStudy (allank):

another question: are you allowed to use a graphing calc?

OpenStudy (anonymous):

yes

OpenStudy (allank):

Without the calc, I believe plugging in the choices in the question (-x-4<=y<=x-5) will work.

OpenStudy (anonymous):

thanks

OpenStudy (allank):

With the graphing calc, the process is a bit longer, but worth knowing. Graph the the two functions: y= x - 5 and y=-x - 4 The two linear graphs should intersect at some point. Sketch out the graphs. For each graph, 1. pick a point on one side of the graph and plug in the x and y values at that point, into the original inequality. 2. If the result is a true inequality, shade the other side of the graph to indicate it is the unwanted region. If the resulting inequality is untrue, shade that side of the graph. Eventually, you should have an unshaded region bound by the two graphs. Then check the choices given. The correct one is that which is located in the unshaded region.